OpenStudy (anonymous):

Evaluate the log: log(base:subscript 7)of (1/49)

OpenStudy (anonymous):

-2

OpenStudy (amistre64):

log7(1/49) ; division in logs is subtraction; since they are exponents at heart log7(1) - log7(49)

OpenStudy (amistre64):

log7(1) = 0 why? because 7^0 = 1 log7(49) = 2 why? because 7^2 = 49 0 - 2 = -2